# VinciCoder: Unifying Multimodal Code Generation via Coarse-to-fine Visual Reinforcement Learning

### v1

@@ Abstract
Multimodal code generation has garnered significant interest within the research community. Despite the notable success of recent vision-language models (VLMs) on specialized tasks like Chart-to-code generation, their reliance on single-task training regimens fosters a narrow paradigm that hinders the development of generalized VIsioN Code Intelligence. In this work, we introduce VinciCoder, a unified multimodal code generation model that addresses this limitation via a two-stage training framework. We begin by constructing a large-scale Supervised Finetuning (SFT) corpus comprising 1.6M image-code pairs for tasks involving direct code generation and visual-based code refinement. Subsequently, we introduce a Visual Reinforcement Learning (ViRL) strategy, which employs a coarse-to-fine reward mechanism to improve visual fidelity by calculating visual similarity across local and global image patches. Extensive experiments on various multimodal code generation benchmarks demonstrate that VinciCoder achieves state-of-the-art performance, underscoring the effectiveness of our coarse-to-fine ViRL strategy. The code and model will be available at https://github.com/DocTron-hub/VinciCoder.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have led to significant breakthroughs in code generation. Leading models, such as Gemini-2.5 [10] and Qwen3-Coder [44], now demonstrate the powerful ability to follow complex textual instructions and generate executable code across multiple languages like Python and Java. Beyond purely textual descriptions, a growing body of research begins to explore multimodal code generation, focusing on visual inputs like charts and web screenshots [38, 54, 17]. These visual modalities are inherently more information-dense than natural language, presenting a more complex challenge for model comprehension and code synthesis.
@@
However, in the multimodal domain, existing works generally focus on training task-specific vision-language models (VLMs). For instance, ChartCoder [54] excels at chart-to-code generation utilizing 160k training samples, while Web2Code [50] constructs a large-scale training corpus for the web-to-HTML domain. Although these specialized models achieve state-of-the-art performance on their respective tasks, their narrow training scopes fundamentally limit their ability to generalize into a unified multimodal code generation framework. Thus, developing a unified multimodal code generation model [16, 35], powered by recent advances in foundation models and reinforcement learning (RL), has emerged as a significant and challenging research direction.
@@
To address these challenges, we introduce VinciCoder, a unified Vision-Language Model (VLM) for multimodal code generation trained via a two-stage SFT-ViRL strategy. For the SFT stage, we construct a large-scale corpus containing 1.3M direct generation samples and 300k examples for a novel task we introduce as code refinement. Distinct from conventional instruction-based editing, this task conditions the model on a target image and a deliberately flawed code snippet that contains logical errors or yields only partial renderings. The model’s objective is to rectify this implementation, generating code whose visual output precisely aligns with the target image. After the SFT phase, we employ RL to further optimize executability and visual fidelity of the generated code. Conventional methods like Reinforcement Learning with Verifiable Rewards (RLVR) are intractable for this domain, as they rely on rule-based textual rewards that are hard to generalize. To overcome this limitation, we propose Visual Reinforcement Learning (ViRL), a framework that pivots the reward mechanism from the textual to the visual domain. ViRL replaces brittle, rule-based validation with a direct reward signal based on perceptual similarity, ensuring scalability and language-agnosticism for the multimodal code generation task.
@@
Furthermore, to robustly compare high-resolution images, our ViRL framework introduces a coarse-to-fine reward function that simultaneously evaluates global structure from downsampled views and local details from segmented image patches. Our experimental results validate this two-stage approach: VinciCoder-SFT, trained on our curated corpus, already establishes a strong baseline, outperforming prior methods on various multimodal code generation benchmarks. The subsequent application of our ViRL strategy then elevates this performance to state-of-the-art (SOTA) levels. To our knowledge, VinciCoder is the first VLM to apply RL for domain-agnostic visual fidelity enhancement of generated code. In summary, our contributions are as follows:
  • We propose VinciCoder, a unified VLM that leverages a two-stage SFT-RL training strategy and visual-based refinement tasks to specialize in diverse multimodal code generation tasks.
@@
  • We conduct a comprehensive evaluation of VinciCoder across diverse multimodal code generation tasks. The results demonstrate that VinciCoder outperforms existing open-source approaches across benchmarks.
@@ 2.1 MLLMs for Code Generation
Multimodal code generation has received considerable attention, with notable advancements in generating code for visual artefacts such as charts, webpages, scalable vector graphics (SVG), and scientific plots. In the chart domain, previous works [54, 52] focus on constructing large-scale corpora to facilitate SFT. Subsequent works [5, 37] integrate RL to further the fidelity of the generated code. In the web domain, data-centric approaches range from generating synthetic pairs with LLMs [50, 19] to collecting massive real-world datasets [12]. Beyond data construction, recent works like LayoutCoder [42] and LatCoder [13] incorporate layout-aware frameworks to produce more structurally accurate HTML code. Unlike chart and web domains, SVG code offers a rich source of paired data for real-world icons and illustrations. StarVector [28] and OmniSVG [47] leverage this to construct datasets for training on image-to-SVG and text-to-SVG tasks. Similarly, numerous real-world scientific plots and corresponding code are utilized for training. MathCoder-VL [39] leverages the real-world Datikz dataset [4, 3] to train an image-to-code model capable of iteratively converting scientific diagrams into LaTeX code.
However, these methods are typically task-specific, constrained to homogeneous visual patterns and singular code languages. While recent research has shifted towards unified models [16, 35], these efforts rely almost exclusively on SFT, which is insufficient for ensuring either code executability and visual fidelity.
@@ 2.2 Multimodal Code Generation Benchmarks
Besides constructing code MLLMs, many multimodal code generation benchmarks have been proposed for evaluation. Previous works generally focus on evaluating direct generation capacity within task-specific domains. For instance, benchmarks such as ChartMimic [45] and Plot2Code [41] evaluate chart-to-code generation capabilities. Other works, including Design2Code [34], UniSVG [21], and Image2Struct [27], assess the code generation of corresponding visual inputs. Also, some benchmarks evaluate beyond direct generation, including interaction [43, 22] and editing generation [7, 53, 46]. Recently, with the growing capacities of VLMs, many new and complex benchmarks have been introduced. Artifactsbench [51] and DCG-Bench [20] propose diverse webpage and chart code generation tasks with dynamic visual images and code complexity.
@@ 2.3 RL for MLLM
Inspired by the success of Group Relative Policy Optimization (GRPO) in post-training LLMs [31, 14], RL has garnered significant attention from the research community. Recently, a growing body of work has focused on applying RL to MLLMs to enhance their vision-language reasoning capabilities. Pioneering efforts, such as Vision-R1 [15], VLM-R1 [32], and R1-OneVision [48], first utilize Chain-of-Thought (CoT) data for SFT to establish a reasoning baseline and employ RL to advance the model policy towards generating coherent answers. This RL framework has proven highly versatile, with successful adaptations to a range of fundamental visual tasks, such as grounding [56, 24] and question answering (QA) [36, 6] and segmentation [23]. Besides constructing rewards from generated text contents, another research direction involves leveraging visual feedback to formulate reward functions. For instance, RVRL [8] and RLRF [29] utilize MLLMs and pretrained Vision Transformers (ViTs) to score generated images, thereby providing a reward signal for RL training.
@@ 3.2 Data Construction
To construct large-scale multimodal code generation data, we curate datasets from various open-source datasets and generate missing image-code types. All code is re-executed to render the corresponding images. We divide according to image types for illustration.
@@ 3.2.1 SFT Data
Chart-to-code. For our chart-to-code task, we curate training data from MCD [16] and MSRL [5]. Our process begins with preprocessing the Chart2Code-160k subset of MCD, where we correct syntax errors using the Ruff linter and custom heuristics. Simultaneously, we curate a diverse 100k-sample subset from the MSRL data using perceptual hashing (pHash) and Mini-Batch K-Means clustering. To construct the code refinement dataset, we first train a specialized chart-to-code VLM on previously curated data. This model is then employed to generate initial code for a new, non-overlapping set of 100k MSRL samples. Acknowledging that the model’s first-pass generations are often imperfect, we form the final refinement dataset by pairing this generated code with the corresponding images and ground-truth targets.
Web-to-HTML. For the web-to-HTML task, we curate training data from MCD [16], Web2M [12], and Web2Code [50]. We first preprocess the Web2M collection, filtering for English-language entries and removing samples with hyperlinks or embedded images. This procedure results in a curated dataset of 60k entries. Following a pipeline similar to our chart-to-code process, we then train a dedicated web-to-code VLM using the combined data from MCD and the filtered Web2M set. Subsequently, this model generates preliminary code for the sampled 100k instances from the Web2Code dataset, forming the final refinement set.
Image-to-SVG. For the image-to-SVG task, our primary training corpus is the ISVGEN subset from UniSVG [21], which contains 360k data pairs. We also construct a refinement dataset with different strategies from the chart and web tasks. Specifically, after training a dedicated VLM on the ISVGEN data, we employ this model to generate code drafts for 100k samples sampled from the ISVGEN training subset itself to constitute the refinement dataset. This strategy is motivated by the observed training instability. The high, fluctuating final training loss suggests the model has not yet memorized the training set, thus validating these samples as non-trivial targets for refinement.
Image-to-Latex. For the image-to-LaTeX task, we curate data from the DaTikZ-v3 [3] and Cosyn-400k [49]. To ensure consistent and valid outputs, we first standardize all code by encapsulating it within a standalone TikZ environment. This step is designed to produce tightly-cropped figures, preventing the generation of full A4-sized PDF pages. Then, we re-execute each sample to validate its integrity, filtering out any instances that result in rendering errors or produce multi-page PDF outputs.
@@ 3.2.2 RL Data
For the RL phase, we construct a new dataset spanning five distinct domains, ensuring it is mutually exclusive with our SFT data. The curation process for each domain is as follows: (i) Chart-to-code: We directly utilize the 11k image-code pairs from the second RL-stage subset of the MSRL dataset [5]. (ii) Web-to-HTML: We sample 9k examples from Web2Code [50] and employ Gemini-2.5-Flash [10] to refine the HTML code. This is achieved by prompting the model with varied reference tags and diverse instructions to enhance the visual complexity and quality of the generated webpages. (iii) Image-to-SVG: Using our trained SFT model as a quality filter, we curate the SVG-icons dataset [28], retaining 10k samples that are both high-fidelity and challenging for the model to generate. (iv) Image-to-LaTeX: Leveraging the recently published ImgCode-8.6M dataset [39], we draw samples from its arxiv-woc-680k category. To rectify observed issues such as inconsistent output scaling and low content diversity, we implemented a hybrid filtering pipeline combining both rule-based and model-based approaches, yielding a final, curated set of 10k examples. (v) Chemical Images: We synthesize a new collection of 2k molecular images by rendering unique SMILES strings with both RDKit [18] and Indigo [11].
@@
Scientific Plots Multi Types 328k 2k
@@ 3.3.2 Reinforcement Learning (RL)
To address these limitations, we propose ViRL, a reinforcement learning strategy designed to optimize for visual fidelity. We first render the generated code into an image and then employ the self-supervised ViT, DINOv2 [26], to encode both the rendered and the input source images into visual embeddings. The reward function is subsequently defined as the normalized cosine similarity between these two visual embeddings.
@@
Visual Reinforcement Learning (ViRL) During the ViRL stage, we employ the Group Relative Policy Optimization (GRPO) [31] algorithm to fine-tune the model. A primary advantage of GRPO is its independence from a separate critic model, a key component in methods like PPO [30]. For a given input query $x$, the algorithm first samples a set of $G$ responses, $\{o_{1},o_{2},\dots,o_{G}\}$, from the current policy $\pi_{old}$. Each response receives a reward $R_{i}$ and the group-normalized advantage for the $i$-th response at time step $t$ is:
@@
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid q,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid q,o_{i,<t})}.$ (7)
The reward $R$ is the combination of coarse-to-fine visual reward $R_{v}$ and language alignment reward $R_{l}$, which is calculated as
$R_{i}=\omega_{v}\frac{1}{N}\sum_{i=1}^{N}R_{i,v}+\omega_{l}R_{l}.$ (8)
When the generated code fails to render images, the $R_{v}$ will be set to 0. The $\omega_{v}$ and $\omega_{l}$ are set to 0.9 and 0.1, respectively. The framework of our proposed coarse-to-fine reward mechanism and ViRL strategy are denoted in Figure 3.

### v2

@@ Abstract
Multimodal code generation has garnered significant interest within the research community. Despite the notable success of recent vision-language models (VLMs) on specialized tasks like chart-to-code generation, their reliance on single-task training regimens fosters a narrow paradigm that hinders the development of generalized VIsioN Code Intelligence. In this work, we introduce VinciCoder, a unified multimodal code generation model that addresses this limitation via a two-stage training framework. We begin by constructing a large-scale Supervised Finetuning (SFT) corpus comprising 1.6M image-code pairs for tasks involving direct code generation and visual-based code refinement. Subsequently, we introduce a Visual Reinforcement Learning (ViRL) strategy, which employs a coarse-to-fine reward mechanism to improve visual fidelity by calculating visual similarity across local and global image patches. Extensive experiments on diverse multimodal code generation benchmarks demonstrate that VinciCoder achieves state-of-the-art performance, surpassing recent open-source models. The ablation study further validates the effectiveness of our proposed coarse-to-fine ViRL strategy. The data, code and model is available at https://github.com/DocTron-hub/VinciCoder.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have led to significant breakthroughs in code generation. Leading models, such as Gemini-2.5 [comanici2025gemini] and Qwen3-Coder [yang2025qwen3], demonstrate powerful abilities to follow complex textual instructions and generate executable code across multiple programming languages like Python and Java. Beyond purely textual descriptions, a growing body of research begins to explore multimodal code generation, focusing on visual inputs like charts and web screenshots [wan2024automatically, jiang2025screencoder]. Visual modalities are inherently more information-dense than natural language, presenting a more complex challenge for model comprehension and code synthesis.
@@
However, in the multimodal domain, existing works generally focus on training task-specific vision-language models (VLMs). For instance, ChartCoder [zhao2025chartcoder] excels at chart-to-code generation utilizing 160k training samples, while Web2Code [yun2024web2code] constructs a large-scale training corpus for the web-to-HTML domain. Although these specialized models achieve superior performance on their respective tasks, their narrow training scopes and foundation in purely supervised fine-tuning (SFT) fundamentally limit their ability to generalize into a unified multimodal code generation framework. Developing a unified multimodal code generation model [jiang2025viscodex, sun2025januscoder], powered by recent advances in foundation models and reinforcement learning (RL), has emerged as a significant and challenging research direction.
@@
To address these challenges, we introduce VinciCoder, a unified vision-language model (VLM) for multimodal code generation trained via a two-stage SFT-ViRL strategy. For the SFT stage, we construct a large-scale corpus containing 1.3 million direct generation data and 300k samples for a novel visual-based code refinement task. Distinct from conventional instruction-based editing, this task conditions the model on a target image and a flawed code snippet that contains logical errors or yields only partial renderings. The objective is for the model to rectify this implementation, generating code whose visual output precisely aligns with the target image. After the SFT phase, we propose a visual reinforcement learning (ViRL) strategy to further optimize executability and visual fidelity of the generated code. Previous reinforcement learning with verifiable reward (RLVR) methods are intractable for code generation tasks, due to their brittle rule-based textual rewards. To overcome this limitation, we propose the ViRL, a framework that pivots the reward mechanism from the textual to the visual domain. Instead of brittle, rule-based validation, ViRL leverages a direct reward signal from perceptual similarity, ensuring scalability and language agnosticism for the multimodal code generation task.
@@
Furthermore, to robustly compare high-resolution images, our ViRL framework introduces a coarse-to-fine reward function that simultaneously evaluates global structure from downsampled views and local details from segmented image patches. We conduct extensive experiments on diverse multimodal code generation benchmarks, benchmarking our approach against leading contemporary models. The results demonstrate the efficacy of our two-stage strategy. First, VinciCoder-SFT, trained on our curated corpus, already establishes a strong baseline by surpassing existing models. The application of our ViRL strategy further enhances the performance, setting a new state-of-the-art (SOTA) across these benchmarks. To our knowledge, VinciCoder is the first unified model to leverage RL for domain-agnostic visual fidelity in multimodal code generation. In summary, our contributions are as follows:
  • We propose VinciCoder, a unified multimodal code generation model trained with a two-stage SFT-RL strategy. The SFT phase leverages a large-scale corpus featuring a novel visual-based refinement task, building a robust, generalizable foundation.
@@
  • We conduct a comprehensive evaluation of VinciCoder across diverse multimodal code generation tasks. The results demonstrate that VinciCoder outperforms recent open-source approaches across benchmarks.
@@ 2.1 MLLMs for Code Generation
Multimodal code generation has received considerable attention, with notable advancements in generating code for visual artefacts such as charts, webpages, scalable vector graphics (SVG), and scientific plots. In the chart domain, previous works [zhao2025chartcoder, zhang2025boosting] focus on constructing large-scale corpora to facilitate SFT. Subsequent works [chen2025breaking, tan2025chartmaster] integrate RL to further the fidelity of the generated code. In the web domain, data-centric approaches range from generating synthetic pairs with LLMs [yun2024web2code, laurenccon2024unlocking] to collecting massive real-world datasets [gui2025webcode2m]. Beyond data construction, recent works like LayoutCoder [wu2025mllm] and LatCoder [gui2025latcoder] incorporate layout-aware frameworks to produce more structurally accurate HTML code. Unlike chart and web domains, SVG code offers a rich source of paired data for real-world icons and illustrations. StarVector [rodriguez2025starvector] and OmniSVG [yang2025omnisvg] leverage this to construct datasets for training on image-to-SVG and text-to-SVG tasks. Similarly, numerous real-world scientific plots and corresponding code are utilized for training. MathCoder-VL [wang2025mathcoder] leverages the real-world Datikz dataset [belouadi2025tikzero, belouadi2024detikzify] to train an image-to-code model capable of iteratively converting scientific diagrams into LaTeX code.
However, these methods are typically task-specific, constrained to homogeneous visual patterns and singular code languages. While recent research has shifted towards unified models [jiang2025viscodex, sun2025januscoder], these efforts rely almost exclusively on SFT, which is insufficient for ensuring either code executability and visual fidelity.
@@ 2.2 Multimodal Code Generation Benchmarks
Besides constructing code MLLMs, many multimodal code generation benchmarks have been proposed for evaluation. Previous works generally focus on evaluating direct generation capacity within task-specific domains. For instance, benchmarks such as ChartMimic [yang2024chartmimic] and Plot2Code [wu2024plot2code] evaluate chart-to-code generation capabilities. Other works, including Design2Code [si2024design2code], UniSVG [li2025unisvg], and Image2Struct [roberts2024image2struct], assess the code generation of corresponding visual inputs. Also, some benchmarks evaluate beyond direct generation, including interaction [xiao2024interaction2code, li2024sketch2code] and editing generation [chen2025svgenius, zhao2025chartedit, yang2025chartm]. Recently, with the growing capacities of VLMs, many new and complex benchmarks have been introduced. Artifactsbench [zhang2025artifactsbench] and DCG-Bench [li2025opusanimation] propose diverse webpage and chart code generation tasks with dynamic visual images and code complexity.
@@ 2.3 RL for MLLM
Inspired by the success of Group Relative Policy Optimization (GRPO) in post-training LLMs [shao2024deepseekmath, guo2025deepseek], RL has garnered significant attention from the research community. Recently, a growing body of work has focused on applying RL to MLLMs to enhance their vision-language reasoning capabilities. Pioneering efforts, such as Vision-R1 [huang2025vision], VLM-R1 [shen2025vlm], and R1-OneVision [yang2025r1], first utilize Chain-of-Thought (CoT) data for SFT to establish a reasoning baseline and employ RL to advance the model policy towards generating coherent answers. This RL framework has proven highly versatile, with successful adaptations to a range of fundamental visual tasks, such as grounding [zhou2025gui, liu2025visual] and question answering (QA) [tan2025reason, chen2025chart] and segmentation [liu2025seg]. Besides constructing rewards from generated text contents, another research direction involves leveraging visual feedback to formulate reward functions. For instance, RRVF [chen2025learning] and RLRF [rodriguez2025rendering] utilize MLLMs and pretrained Vision Transformers (ViTs) to score generated images, thereby providing a reward signal for RL training.
@@ 3.2 Data Construction
To construct large-scale multimodal code generation data, we curate datasets from various open-source datasets and generate missing image-code types. All code is re-executed to render the corresponding images.
@@ 3.2.1 SFT Data
Chart-to-code. For our chart-to-code task, we curate training data from MCD [jiang2025viscodex] and MSRL [chen2025breaking]. Our process begins with preprocessing the Chart2Code-160k subset of MCD, where we correct syntax errors using the Ruff linter and custom heuristics. Also, we curate a diverse 100k-sample subset from the MSRL data using perceptual hashing (pHash) and mini-batch K-means clustering. To construct the code refinement dataset, we first train a specialized chart-to-code VLM on previously curated data. This model is then employed to generate initial code for a new, non-overlapping set of 100k MSRL samples. Recognizing that first-pass generations are often imperfect, our refinement dataset pairs this generated code with its target image and ground-truth. The format is illustrated in the Appendix.
Web-to-HTML. For the web-to-HTML task, we curate training data from MCD [jiang2025viscodex], Web2M [gui2025webcode2m], and Web2Code [yun2024web2code]. We first preprocess the Web2M collection, filtering for English-language entries and removing samples with hyperlinks or embedded images. This procedure results in a curated dataset of 60k entries. Following a pipeline similar to our chart-to-code process, we then train a dedicated web-to-code VLM using the combined data from MCD and the filtered Web2M set. Subsequently, this model generates preliminary code for the sampled 100k instances from the Web2Code dataset, forming the final refinement set.
Image-to-SVG. For the image-to-SVG task, our primary training corpus comes from the UniSVG ISVGEN subset[li2025unisvg], which contains 360k data pairs. We also construct a refinement dataset with different strategies from the chart and web tasks. Specifically, after training a dedicated VLM on the ISVGEN data, we employ this model to generate code drafts for 100k samples sampled from the ISVGEN training subset itself to construct the refinement dataset. This strategy is motivated by the observed training instability. The high, fluctuating final training loss suggests the model has not yet memorized the training set, thus validating these samples as non-trivial targets for refinement.
Image-to-Latex. For the image-to-LaTeX task, we curate data from the DaTikZ-v3 [belouadi2024detikzify] and Cosyn-400k [yang2025scaling]. To ensure consistent and valid outputs, we first standardize all code by encapsulating it within a standalone TikZ environment. This step is designed to produce tightly-cropped figures, preventing the generation of full A4-sized PDF pages. Then, we re-execute each sample to validate its integrity, filtering out any instances that result in rendering errors or produce multi-page PDF outputs.
@@ 3.2.2 RL Data
For the RL phase, we construct a new dataset spanning five distinct domains, ensuring it is mutually exclusive with our SFT data. The curation process for each domain is as follows: (i) Chart-to-code: We directly utilize the 11k image-code pairs from the second RL-stage subset of the MSRL dataset [chen2025breaking]. (ii) Web-to-HTML: We sample 9k examples from Web2Code [yun2024web2code] and employ Gemini-2.5-Flash [comanici2025gemini] to refine the HTML code. This is achieved by prompting the model with varied reference tags and diverse instructions to enhance the visual complexity and quality of the generated webpages. (iii) Image-to-SVG: Using our trained SFT model as a quality filter, we curate the SVG-icons dataset [rodriguez2025starvector], retaining 10k samples that are both high-fidelity and challenging for the model to generate. (iv) Image-to-LaTeX: Leveraging the recently published ImgCode-8.6M dataset [wang2025mathcoder], we draw samples from its arxiv-woc-680k category. To rectify observed issues such as inconsistent output scaling and low content diversity, we implemented a hybrid filtering pipeline combining both rule-based and model-based approaches, yielding a final, curated set of 10k examples. (v) Chemical Images: We synthesize a new collection of 2k molecular images by rendering unique SMILES strings with both RDKit [landrum2013rdkit] and Indigo [IndigoToolkit].
@@
Scientific Plots Multiple 238k 2k
@@ 3.3.2 Reinforcement Learning (RL)
To address these limitations, we propose ViRL, a reinforcement learning strategy designed to optimize for visual fidelity. We first render the generated code into an image and then employ the self-supervised ViT, DINOv2 [oquab2023dinov2], to encode both the rendered and the input source images into visual embeddings. The reward function is subsequently defined as the normalized cosine similarity between these two visual embeddings.
@@
Visual Reinforcement Learning (ViRL) During the ViRL stage, we employ the Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] algorithm to fine-tune the model. A primary advantage of GRPO is its independence from a separate critic model, a key component in methods like PPO [schulman2017proximal]. For a given input query $x$, the algorithm first samples a set of $G$ responses, $\{o_{1},o_{2},\dots,o_{G}\}$, from the current policy $\pi_{old}$. Each response receives a reward $R_{i}$ and the group-normalized advantage for the $i$-th response at time step $t$ is:
@@
$r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}\mid x,o_{i,<t})}{\pi_{\theta_{\text{old}}}(o_{i,t}\mid x,o_{i,<t})}.$ (7)
The reward $R$ is the combination of multiple coarse-to-fine visual rewards $R_{v}$ and language alignment reward $R_{l}$, which is calculated as
$R=\omega_{v}\frac{1}{N}\sum_{i=1}^{N}R_{i,v}+\omega_{l}R_{l}.$ (8)
When the generated code fails to render images, the $R_{v}$ will be set to 0. The $\omega_{v}$ and $\omega_{l}$ are set to 0.9 and 0.1, respectively. Figure 3 shows the framework of our proposed coarse-to-fine reward mechanism and ViRL strategy.